# Safe MPC Alignment with Human Directional Feedback

### v1

@@ VII-B Participants, Procedures, and Metrics
All participants recruited for the two simulation games in Section VI-D also take part in the hardware experiment user study. They are told of the objective before starting trials. Similar to part VI-C, one episode is defined as one robot motion from the starting gripper position to the target position. One trial means one run of the control with multiple episodes (with a maximum of 15 corrections). During corrections, the nozzle of the water pot is sealed, and we switch to the unsealed nozzle and measure the amount of water poured into the flower pot after the user is satisfied with the motion of the robot motion. The success trial is defined when the amount of poured water is over 50g. Each user performs 5 trials, and the same metrics in Section VI-D are recorded.
VII-C Result and Analysis
@@
In real-world robot experiments, the success rate is $84.0\%\pm 16.7\%$ and the average number of corrections in successful trials is $6.88\pm 3.26$. The results show the efficacy and efficiency of the proposed method in real-world robot learning applications. To illustrate the robot learning progress, in Fig. 19, we use one trial of the user study to show the robot behavior at different numbers of human corrections.
@@
Since the task of this real-world Franka arm experiment is similar to the simulated Franka arm reaching game in Section VI-C, we compare the user study results for both and have the following conclusion and analysis.
@@
(I) The real-world experiment has a slightly lower success rate: While the difference between simulations and real-world experiments is small in success rate, we have identified some factors during our experiment that may contribute to such lower success rate, which includes both the human factors and hardware factors. On human factors, most of the participants have no experience or knowledge in operating robot hardware, and thus when directly interacting with hardware, their corrections can be conservative or ineffective due to safety considerations. On the hardware side, some error sources, such as noisy correction signals, and the possible timing lag between the recorded correction-happening robot state and the human intended correction-happening state, can contribute to the failure of the learning.
@@
(II) The real-world experiment has a less number of human corrections: This could be attributed to two reasons. First, the constraint needed for accomplishing the water pouring task in the real-world experiment could be more relaxed than the constraint for accomplishing the narrow slot passing task in the simulation game in Section VI-C. This makes the safety constraints easy to learn for the real-world task. Second, different from simulation games, where human users can only perceive the motion of the robot arm from the screen, a human user has a better perception of the robot’s motion in a real-world experiment. A good perception of the robot’s motion could help humans give more effective feedback each time, reducing the total number of human interactions for successful robot learning.
VII-D Discussion
@@
Based on the observations in both the hardware and simulation user studies, we discuss some design choices and potential limitations of the proposed safe MPC alignment algorithm.
VII-D1 Learning Task-Independent Safety Constraints
@@
In the real-world and simulation experiments, due to the insufficiency of task sampling, the learned safety constraint can be task-dependent, instead of task-independent. For instance, in the drone navigation game in Section VI-B, since the three task targets are placed above the gate, the learned constraint shown in Fig. 15 can only reflect the upper edge of the gate constraints. One limitation of this is that the learned task constraints have poor generalization to new tasks, which are far from previously encountered. For example, in drone navigation, if a new target is placed below the gate, the drone with the learned safety constraint Fig. 15 will fail.
@@
To improve the generalization of the learned safety constraints or learning task-independent safety constraints, one solution is to increase the diversity of task samples. By doing so, it is expected that the robot can receive human correction on the entire boundary of the safety region, making the learned safety constraints task irrelevant.
VII-D2 Robustness Against Human Irrationality
@@
Our proposed safe MPC alignment method is based on the cutting hypothesis. In our formulation, we define a hypothesis cut (7) based on an average of multiple human corrections, such an averaging treatment is to mitigate human mistakes for one cut. But still, in practice, human irrationality can lead to the wrong cut of the hypothesis space, which constitutes the main reason for failure in user studies. For example, in the drone navigation game, we found that failure game trials usually happen when a human gives too many wrong directional corrections at the beginning of the trial. A failure case is shown in Fig. 20.
@@
Figure 20: One illustration of the failure case in the drone user study. A human user makes a false "up" correction at the beginning and fails to correct the motion of the drone.
@@
To further improve the robustness of the method against human mistakes, two extensions of the current method can be used on top of our proposed algorithm. The first is to add a relaxation factor $\epsilon\geq 0$ in our cutting hypothesis. For example, we can re-write (7) as
@@
Another extension is to add a selection mechanism on human correction. For example, one can use some statistical method to distinguish and rule out mistake corrections before performing a hypothesis cut. We will leave those possible extensions in our future work.
VIII Conclusion
This paper proposes safe MPC alignment method to online learn a safety constraint in robot safe MPC controllers from human directional feedback. The method, named hypothesis space cutting, has geometric interpretation: it exponentially cuts the hypothesis space based on each received human directional feedback. The method has a provable guarantee for the upper bound of the total number of human feedback, and also can declare hypothesis misspecification. A numeric example, two user studies on simulation games, and a user study on the real-world robot arm, covering varying robot tasks, were used to validate the effectiveness and efficiency of the proposed approach. The results show that the proposed method can learn safety constraints effectively and efficiently. To our knowledge, this is the first certifiable method to enable online learning of safety constraints with human feedback.
@@ A-B Proof of Lemma 4
For any $i=1,2,3,...$, the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}=\{\boldsymbol{u}^{\boldsymbol{\theta}_{i},\gamma}_{0:T-1},\boldsymbol{x}^{\boldsymbol{\theta}_{i},\gamma}_{0:T}\}$ is a solution to the optimization (3), with $\boldsymbol{\theta}=\boldsymbol{\theta}_{i}$. Recall Lemma 1, $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$. Extending the feature-weight expression of $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$, one can get $\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$. Since the optimal solution $\boldsymbol{u}^{\boldsymbol{\theta}_{i},\gamma}_{0:T-1}$ is a minimizer to (3), it means
@@ A-C Proof of Lemma 5
$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r}.$ (49)

### v2

@@ VIII-B Participants, Procedures, and Metrics
All participants recruited for the previous two simulation games in Section VII-B also participate the hardware experiment. They are told of the objective before starting trials. One episode is defined as one robot motion from the starting position to the targets, during which the human can apply corrections at any time instances. One learning trial may include multiple episodes (with a maximum of 15 corrections). During learning, the nozzle of the water pot is sealed, and we switch to the unsealed nozzle and measure the amount of water poured into the flower pot after the user is satisfied with the motion of the robot motion. The success learning trial is defined when the amount of poured water is over 50g. Each user performs 5 trials, and the same metrics in Section VII-B are recorded.
VIII-C Result and Analysis
@@
In the real-world robot experiment, the success rate is $84.0\%\pm 16.7\%$ and the average number of corrections needed for successful trials is $6.88\pm 3.26$. The results show the efficacy and efficiency of our method in real-world robot learning applications. To illustrate the robot learning progress, in Fig. 22, we use one trial to show the robot behavior at different numbers of human corrections.
@@
Since this real-world Franka arm experiment is similar to the simulated Franka arm reaching game in Appendix B-D, we compare the user study results for both and have the following conclusion and analysis.
@@
(I) The real-world experiment has a slightly lower success rate: While the difference between simulations and real-world experiments is small in success rate, we have identified some factors during our experiment that may contribute to such lower success rate, including both the human aspects and hardware aspects. On human aspects, most of the participants have no experience or knowledge in operating robot hardware, and thus when directly interacting with hardware, their corrections can be conservative or ineffective due to safety considerations. On hardware side, noisy correction signals and the possible timing delay between the recorded human correction and robot state can contribute to the failure of the learning.
@@
(II) The real-world experiment has a less number of human corrections: This could be attributed to two reasons. First, the constraint needed for accomplishing the water pouring task could be more flexible than that for reaching task in the simulation game Appendix B-D. This makes the safety constraints easy to learn for the real-world task. Second, different from simulation games, where human users can only perceive the robot arm from the screen, a human user has a better perception of the robot’s motion in real world. A good perception of the robot’s state could help humans give more effective feedback each time, reducing the total number of human interactions for successful robot learning.
VIII-D Discussion
@@
Based on the observations in both the hardware and simulation user studies, we discuss some design choices and potential limitations of the proposed method.
VIII-D1 Learning Task-Independent Safety Constraints
@@
In the real-world and simulation experiments, due to the insufficiency of task sampling, the learned safety constraint could be task-specific, instead of task-independent. For instance, in the drone navigation game, since the three task targets are placed above the gate, the learned constraint shown in Fig. 19 can only reflect the upper edge of the unknown gate. One limitation of this is that the learned task constraints have poor generalization to new tasks. For example, in drone navigation, if a new target is placed below the gate, the drone with the learned safety constraint Fig. 19 will fail.
@@
To improve the generalization of the learned safety constraints or learning task-independent safety constraints, one solution is to increase the diversity of task samples. By doing so, it is expected that the robot can receive human correction on the entire boundary of the safety region, making the learned safety constraints reusable in new tasks.
VIII-D2 Robustness Against Human Mistakes
@@
The proposed safe MPC alignment is based on the hypothesis cutting. In our formulation, a hypothesis cut (7) is based on an average of multiple human corrections, such an averaging treatment can mitigate human mistakes for one cut. But still, in practice, human irrationality can lead to the wrong cut of the hypothesis space, which is the main reason for failure in user studies. For example, in the drone navigation game, we found that failure game trials usually happen when a human gives too many wrong directional corrections. A failure case is in Fig. 23.
@@
Figure 23: One illustration of the failure case in the drone game user study. A human user makes a false "up" correction at the beginning and fails to correct the motion of the drone.
@@
To further improve the robustness of the method against human mistakes, two extensions of the current method can be explored. The first is to add a relaxation factor $\epsilon\geq 0$ in our cutting hypothesis. For example, we can re-write (7) as
@@
Another is to add a selection mechanism on human correction. For example, one can use some statistical method to rule out mistake corrections before performing a hypothesis cut. We will leave those possible extensions in our future work.
IX Conclusion
This paper proposes safe MPC alignment method to online learn a safety constraint in safe MPC controllers from human directional feedback. The method, named hypothesis space cutting, has geometric interpretation: it exponentially cuts the hypothesis space based on each received human directional feedback. The method has a provable guarantee for the total number of human feedback needed for learning, and also can declare hypothesis misspecification. Numeric examples, user studies on simulation games, and a user study on the real-world robot arm, covering varying robot tasks, were used to validate the effectiveness of the proposed approach. The results show that the proposed method can learn safety constraints effectively and efficiently. To our knowledge, this is the first certifiable method to enable online learning of safety constraints with human feedback.
@@ A-B Proof of Lemma 3
(14) can be directly obtained from (12), (2) and (5). Based on the definition of the penalty objective function (2), $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is a linear fractional inequality
with a positive denominator $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ due to (12).
Multiplying the denominator on both side and substituting the weight-feature expression of $g_{\boldsymbol{\theta}}$ in (5) yields $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}$ in Lemma 3. The condition $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ itself yields $\ {\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ in Lemma 3.
This completes the proof. ∎
A-C Proof of Lemma 4
Proof.
For any $i=1,2,3,...$, the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ is a solution to the optimization (3), with $\boldsymbol{\theta}=\boldsymbol{\theta}_{i}$. Recall Lemma 1, $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$. Extending the feature-weight expression of $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$, one can get $\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$. Since the optimal solution $\boldsymbol{u}^{\boldsymbol{\theta}_{i},\gamma}_{0:T-1}$ is a minimizer to (3), it means
@@ A-D Proof of Lemma 5
$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r},$ (43)
@@ A-F Proof of Theorem 2
Appendix B Settings of Experiments and User Studies
B-A Inverted Pendulum
In the dynamics of an inverted pendulum (26), the physical parameters is set in Table IV.
TABLE IV: Physics Parameters in Pendulum Numerical Study
Parameter Value Interpretation
$m$ 1kg Weight of the pendulum
$l$ 1m Length of the pendulum
$\mathrm{g}$ $10\mathrm{m/s}^{2}$ Gravitational Acceleration
$d$ 0.1 Dampening Coefficient
B-B Two-link Robot Arm
We assume each links has unit length and mass, and the mass is evenly distributed along the link. The corresponding inertia matrix and Coriolis term can then be calculated.
B-C Drone Navigation Game
B-C1 Environment
The drone starts from $[0.1,y_{\text{init}},0.5]^{\mkern-1.5mu\mathsf{T}}$ with $y_{\text{init}}$ uniformly randomly in an interval $[4,6]$. The starting quaternion $\boldsymbol{q}_{\text{init}}=[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. We set three different targets at positions $[19,1,9]^{\mkern-1.5mu\mathsf{T}}$, $[19,5,9]^{\mkern-1.5mu\mathsf{T}}$ and $[19,9,9]^{\mkern-1.5mu\mathsf{T}}$, with quaternion $[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. For the narrow gate (yellow rectangle) in Fig. 15, both its size and position are unknown, and one only roughly knows its x position $x_{\text{gate}}\approx 9.8$.
where $\beta_{t}$ is a weighting factor to balance the immediate and future constraint satisfaction. If one emphasizes instant safety, more weights will be given to the initial time step. In our implementation, we choose an exponentially decaying $\beta_{t}$ starting at $t=5$, i.e. $\beta_{t}=0,\ t<5$ and $\beta_{t}=0.9^{t-5},\ t\geq 5$ to achieve better predictive safety performance. In (55), each RBF is defined as
Here, the variables $r_{xy,t}=[r_{x,t},r_{y,t}]^{\mkern-1.5mu\mathsf{T}}$ and $r_{xz,t}=[r_{x,t},r_{z,t}]^{\mkern-1.5mu\mathsf{T}}$ with $r_{x},r_{y},r_{z}$ from $\boldsymbol{x}_{t}$. The $i$th center $c_{xy,i}=[c_{x,i},c_{y,i}]^{\mkern-1.5mu\mathsf{T}}$ or $c_{xz,i}=[c_{x,i},c_{z,i}]^{\mkern-1.5mu\mathsf{T}}$ is defined with $c_{x,i}={x_{\text{gate}}}=9.8$, $c_{y,i}$ and $c_{z,i}$ are evenly chosen on interval $[0,10]$. We set $\epsilon=0.45$. The initial hypothesis space $\Theta_{0}$ for possible $\boldsymbol{\theta}$ in (54) is set as $\boldsymbol{c}_{l}=[-80,\dots,-80]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[100,\dots,100]^{\mkern-1.5mu\mathsf{T}}$. $\gamma$ is 50. This initialization blocks the drone from advancing to the gate at the beginning of the game, leading to conservative behavior in the learning process. The drone will proceed toward the gate after few corrections.
B-C2 Human Correction Interface
Human users apply corrections to the drone via a keyboard during the drone flying. The user is informed of the mappings between the key press and the corresponding correction in the game, listed in Table V. The safety constraint (54) in the drone’s MPC policy is immediately updated once a user correction is detected. When the user presses "ENTER", the drone is reset to the initial position and target will randmize.
TABLE V: keyboard commands in uav user study
Key Correction Interpretation
"enter" N/A Reset robot position and switch target.
B-D Franka Constrained Reaching Game
B-D1 Environment
The arm starts at a configuration in which the end-effector position $\boldsymbol{r}_{\text{init}}=[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ and altitude $\boldsymbol{q}_{\text{init}}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The reaching target position $\boldsymbol{r}^{*}$ will be randomly chosen from three options $[-0.6,-0.5,0.4]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.6]^{\mkern-1.5mu\mathsf{T}}$. The reaching target attitude is $\boldsymbol{q}^{*}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The slot size and position are unknown except its rough x-position which is $x_{\text{slot}}\approx-0.15$.
respectively.
The above end-effector dynamics is discretized with $\triangle t=0.1\mathrm{s}$. In (1), the MPC horizon is $T=20$. Let $\boldsymbol{x}^{*}=[\boldsymbol{r}^{*}\ \boldsymbol{q}^{*}]$ denote the target end-effector pose. The step cost $c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})$ and terminal cost $h(\boldsymbol{x}_{T})$ in (1) is set to
Similar to Appendix B-C, to constrain $r_{z}$ and $\boldsymbol{q}$ for proper height and tilt in order to pass through the slot. We use a set of $N=20$ Gaussian radial basis function (RBFs) to parameter the safety constraint:
with
which is used to constrain the $\mathrm{RBF}_{i}$ only taking effect around $x$ range $[x_{\text{slot}}-\frac{w}{2},x_{\text{slot}}+\frac{w}{2}]$ (recall slot’s $x$ location, slot, is a approximate value). In the above definition, $w=0.3$, $k_{z}=0.05$, $\epsilon_{z}=12$ and $\epsilon_{q}=1.8$. The center of positional $\mathrm{RBF}$, ${c_{z,i}}$, is evenly chosen on interval $[0,1]$, and the center of orientation $\mathrm{RBF}$, ${\boldsymbol{q}_{c,i}}$, is generated by $(\cos{\frac{\alpha}{2}},\sin{\frac{\alpha}{2}},0,0)*\boldsymbol{q}_{\text{init}}$ with $\alpha$ evenly choosen from $[-\pi/4,3\pi/4]$ to make sure ${\boldsymbol{q}_{c,i}}$ covers all possible orientations. $\beta_{t}=0.9^{t}$ if $t<5$ and 00 otherwise. The bounds of the initial hypothesis space $\Theta_{0}$ is set to $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[5,\dots,5]^{\mkern-1.5mu\mathsf{T}}$. $\gamma$ is set to 1.5. This initialization leads to a conservative MPC policy, preventing the arm from approaching the target at the beginning.
B-D2 Human Correction Interface
TABLE VI: keyboard commands in robot arm user study
Key Correction Interpretation
"shift" $[0,0,0,-1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go up in world frame
"ctrl" $[0,0,0,1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go down in world frame
"Enter" N/A Reset robot position and switch target.